OpenStudy (anonymous):

i need help with this one 3 - 4x = 10x +10

OpenStudy (anonymous):

-7 = 14x therefore x=-.5

OpenStudy (anonymous):

is this a solution

OpenStudy (anonymous):

Yes -.5 is the correct answer.

OpenStudy (anonymous):

its not giving me that choice on my wrksheet

OpenStudy (anonymous):

What are you choices?

jhonyy9 (jhonyy9):

14x= -7 x= -7/14 x= -1/2

jhonyy9 (jhonyy9):

ok thank you

OpenStudy (anonymous):

x= 120 3 - 4x = 10x + 10 no solution x = 6/7 x = -1